Session: Feb 2004 (session 1)
Wavelength: 6 cm
To give feedback on this experiment, click
Onsala (Michael Lindqvist at Sun Feb 15 08:39:27 2004) Failure :
We observed less than an hour (until around 22:30). Then we got problem with antenna tracking. Problem is still not solved.
Medicina (A.Maccaferri at Sun Feb 15 09:08:39 2004) Minor failures :
Possible fault on tape, check say head 2 not in requested position.
Effelsberg (A. Kraus at Sun Feb 15 12:03:11 2004) Success :
Torun (K.Borkowski at Mon Feb 16 05:23:58 2004) Success :
No known problems.
Jodrell_Bank (Alastair Gunn at Mon Feb 16 12:48:29 2004) Minor failures :
Problem with MERLIN control. On source at 21:50 UT (scan 2).
Vacuum problem occurred around 23:28 UT. Restarted successfully
at 00:14 UT. No other known problems.
Cambridge (Alastair Gunn at Mon Feb 16 12:49:26 2004) Minor failures :
Problem with MERLIN control. On source at 21:52 UT (scan 2).
Phase CAL not present at start - switched on at 22:12 UT. No
other known problems.
Noto (P. Cassaro at Wed Feb 18 11:44:14 2004) Success :
Westerbork (G. Kuper at Wed Feb 25 09:50:30 2004) Minor failures : RECO
Auto bank-switch failed. Done by hand